# Counting points on character varieties

### v1

@@ 2.1.2. Dual root systems
For each root $\alpha\in\Phi$, define the coroot $\check{\alpha}:=\frac{2}{(\alpha,\alpha)}\alpha\in V$. The set of coroots forms the dual root system $\check{\Phi}\subset V$. If $\Psi$ is a subsystem of $\Phi$, then $\check{\Psi}$ is a subsystem of $\check{\Phi}$. However, if $\Psi$ is closed in $\Phi$, then $\check{\Psi}$ need not be closed in $\check{\Phi}$. For instance, $C_{n}\times C_{n}$ is a closed subsystem of $C_{2n}$ but if $n>1$, then $B_{n}\times B_{n}$ is not closed in $B_{2n}$. Indeed, it is easy to see that $B_{n}\times B_{n}$ does not arise in the Borel–de Siebenthal algorithm applied to $B_{2n}$.
@@ 2.1.3. Levi subsystems
A Levi subsystem [^5] In the literature, these are usually called parabolic subsystems, but we find our terminology more useful for comparing with pseudo-Levi and endoscopy subsystems defined below. of $\Phi$ is a subsystem of the form $\Phi\cap E$ where $E\subseteq V$ is a subspace. A Levi subsystem of $\Phi$ is closed in $\Phi$. Moreover, every Levi subsystem of $\Phi$ is of the form $w.\langle S\rangle$, where $S$ is a subset of $\Delta$ and $w\in W$. If $w=1$, then $\langle S\rangle$ is called a standard Levi subsystem. Note that $\Psi$ is a Levi subsystem of $\Phi$ if and only if $\check{\Psi}$ is a Levi subsystem of $\check{\Phi}$.
@@ 2.1.4. Isolated subsystems
A subset $S\subseteq\Phi$ is called isolated in $\Phi$ if $S$ is not contained in a proper Levi subsystem of $\Phi$. One can check that the following are equivalent:
@@ 2.1.5. Pseudo-Levi subsystems
Assume $\Phi$ is irreducible. Let $\theta$ be the highest root of $\Phi$, and $\widetilde{\Delta}:=\Delta\sqcup\{-\theta\}$. A pseudo-Levi subsystem is a subsystem of $\Phi$ of the form $w.\langle S\rangle$, where $w\in W$ and $S$ is a subset of $\widetilde{\Delta}$. If $w=1$, then $\langle S\rangle$ is called a standard pseudo-Levi subsystem. More generally, if $\Phi=\Phi_{1}\sqcup\cdots\sqcup\Phi_{r}$ with $\Phi_{i}$ irreducible, then a subsystem $\Psi\subseteq\Phi$ is said to be a pseudo-Levi if $\Psi_{i}:=\Psi\cap\Phi_{i}$ is a pseudo-Levi inside $\Phi_{i}$ for all $i$. It is clear that every Levi is a pseudo-Levi. The converse is true in type $A$ but false otherwise. For instance, $C_{2}$ has a pseudo-Levi $A_{1}\times A_{1}$ which is not a Levi.
@@ 2.1.6. Isolated pseudo-Levis
Let $\Psi$ be a pseudo-Levi subsystem of an irreducible root system $\Phi$. One can check that the following are equivalent:
@@
    $\Psi=w.\langle S\rangle$ where $S\subset\tilde{\Delta}$ has the same size as $\Delta$.
@@ 2.2.3. Parabolic subgroups
A connected subgroup $P\subseteq G$ is called parabolic if the quotient $G/P$ is proper. Every parabolic subgroup has a Levi decomposition $P=L\ltimes N$, where $N$ is the maximal closed connected normal unipotent subgroup of $P$ and $L$ is a Levi subgroup of $G$. Every Levi subgroup of $G$ arises in this manner. In what follows, we will be using the following facts:
@@ 2.2.4. Pseudo-Levi subgroups
This is a pseudo-Levi subsystem of $\Phi$. Moreover, every pseudo-Levi subsystem arises via the centraliser of some element of $T$, cf. [Deriziotis]. By a theorem of Steinberg, $\pi_{0}(G_{x})$ is isomorphic to a subgroup of $\pi_{1}([G,G])$; thus, if $[G,G]$ is simply connected, then $G_{x}$ is connected [Steinberg75, §2].
@@ 2.2.6. Quasi-isolated elements
A semisimple element $x\in G$ is called quasi-isolated if $G_{x}$ is not in a proper Levi subgroup of $G$. It is called isolated if $G_{x}^{\circ}$ is not in any proper Levi subgroup of $G$. Thus, $x\in T$ is isolated if and only if $\Phi_{x}$ is an isolated (pseudo-Levi) subsystem of $\Phi$. If $x$ is quasi-isolated, then $G_{x}$ is an irreducible subgroup of $G$ (i.e., it does not belong to a proper parabolic). By Schur’s lemma (cf. [Sikora, Proposition 15]) $Z(G_{x})/Z(G)$ is finite.
@@ 3. Generic character varieties
Let $G$ be a connected reductive group over an algebraically closed field $k$. In this section, we define the notion generic tuples of conjugacy classes of $G$ and study the corresponding character varieties. Most of this material appears in [HLRV11] and [Boalch14], but our exposition is slightly different and more suited for our purposes.
@@ 3.1.1. Maximal parabolics
Clearly it is enough to check the condition on maximal proper parabolic subgroups. If $G=\mathrm{GL}_{n}$, this is equivalent to requiring that if a proper non-zero subspace $V\subset k^{n}$ is preserved by $x_{i}\in C_{i}$, $i=1,\ldots,n$, then $\prod_{i=1}^{n}\mathrm{det}(x_{i}|_{V})\neq 1$; thus, we recover the notion of generic in [HLRV11, §2].
@@ 3.1.2. Standard parabolics
It is enough to check the above condition for standard (proper) parabolic subgroups. Here, standard parabolic subgroups are parabolic subgroups containing a chosen Borel subgroup $B\subset G$. Indeed, suppose $P$ is an arbitrary parabolic and choose $g\in G$ such that the conjugate $g.P$ is standard. Then $x_{i}\in P\cap C_{i}$ if and only if $g.x_{i}\in g.P\cap C_{i}$ and $\prod_{i=1}^{n}x_{i}\in[P,P]$ if and only if $g.\prod_{i=1}^{n}x_{i}=\prod_{i=1}^{n}g.x_{i}\in g.[P,P]=[g.P,g.P]$.
@@ 3.1.3. Semisimple part
For each conjugacy class $X$, write $X^{s}$ for the corresponding semisimple class, obtained by taking the semisimple parts of the elements of $X$. Then one can show that $C=(C_{1},\ldots,C_{n})$ is generic if and only if $C^{s}=(C_{1}^{s},\ldots,C_{n}^{s})$ is generic. This is the reductive analogue of the fact that, in $\mathrm{GL}_{n}$, being generic depends only on the eigenvalues.
@@ 3.1.5. Generic elements
We now explain that deciding whether a tuple of (semisimple) conjugacy classes is generic involves checking finitely many conditions.
@@
A tuple $S=(S_{1},\ldots,S_{n})\in T^{n}$ is said to be generic if for every $w=(w_{1},\ldots,w_{n})\in W^{n}$ and every proper standard Levi subgroup $L\subset G$, we have $\prod_{i=1}^{n}w_{i}.S_{i}\notin T\cap[L,L]$.
Let $C_{i}:=G.S_{i}$ be the conjugacy class corresponding to $S_{i}$.
@@
It is clear that if $C$ is generic then so is $S$. For the converse, let $x=(x_{1},\ldots,x_{n})\in C_{1}\times\ldots\times C_{n}$. Let $L\subset G$ be a proper standard Levi subgroup containing $x_{1},\ldots,x_{n}$. We need to show $\prod_{i=1}^{n}x_{i}\notin[L,L]$. We have $x_{i}=l_{i}.t_{i}$ for some $l_{i}\in L$ and $t_{i}\in T$. Now $t_{i}$ and $S_{i}$ are elements of $T$ which are $G$-conjugate; thus, they must actually be conjugate under $W$; i.e., there exists $w_{i}\in W$ such that $t_{i}=w_{i}.S_{i}$. By assumption, $t_{1}\cdots t_{n}\notin[L,L]$ which implies that $x_{1}\cdots x_{n}\notin[L,L]$. Thus, $C$ is generic. ∎
@@ 3.1.6. Existence
The above lemma shows that generic conjugacy classes exist. Indeed, if $L$ is a proper standard levi subgroup, then $T\cap[L,L]$ is a proper closed subvariety of $T$. (At the level of Lie algebras, $\mathfrak{t}=\mathfrak{t}\cap[\mathfrak{l},\mathfrak{l}]\oplus Z(\mathfrak{l})$ for $\mathfrak{t}=\mathrm{Lie}(T)$ and the centre $Z(\mathfrak{l})$ is nontrivial if $\mathfrak{l}$ is a proper Levi subalgebra of $\mathfrak{g}$.) Since there are only finitely many standard Levis $L$ and finitely many $w$’s, we conclude that the generic locus of $T^{n}$ is an open dense subvariety.
@@
Let $\mathbf{T}_{n}\subseteq T^{n}$ denote the subvariety consisting of those tuples $(S_{1},\ldots,S_{n})\in T^{n}$ satisfying $\prod_{i=1}^{n}S_{i}\in[G,G]$. Then the above argument shows that the generic locus of $\mathbf{T}_{n}$ is open dense. Since the (strongly) regular locus is also open dense in $T^{n}$, it follows that conjugacy classes satisfying Assumption 1 exist.
@@ 3.2. Generic character varieties
Let ${r=(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in R}$ and suppose $P$ is a parabolic subgroup of $G$ containing all $a_{i},b_{i},$ and $c_{i}$. Then, $[P,P]$ contains $[a_{1},b_{1}]\cdots[a_{g},b_{g}]$; thus, $[P,P]$ also contains $\prod_{i=1}^{n}c_{i}$. As $C$ is generic, it follows that $P=G$. Thus, $r$ is irreducible and Schur’s lemma (cf. [Sikora, §4]) implies that $\mathrm{Stab}_{G/Z}(r)$ is finite. Note that this group is the same as the stabiliser of $\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle\subseteq G$; thus, by Theorem 7, it is étale.
Stability follows from the fact that whenever a connected reductive group acts on an affine scheme of finite type with finite stabilisers, then every point is stable, cf. [Mumford, Proposition 0.8]. (Alternatively, one can use the Hilbert–Mumford criterion as in [Boalch14].) The statements about $R$ follow from the fact that $1$ is a regular value for $M$, cf. [KNP23, §2] for details. ∎
@@
Next, let ${X}:=R/\!\!/G$ and $[{X}]:=[R/(G/Z)]$ denote, respectively, the generic character variety and character stack associated to $(\Gamma_{g,n},G,C)$. The above theorem implies:
@@ 3.3. Frobenius Mass Formula
We now work over $k=\overline{\mathbb{F}_{q}}$ and assume that $G$ is equipped with an untwisted Frobenius automorphism $F:G\rightarrow G$. We also assume that the conjugacy classes $C=(C_{1},\ldots,C_{n})$ are $F$-stable. Then the action of the Frobenius $F$ on $G$ and $C_{i}$’s, gives an action of $F$ on $R$, ${X}$ and $[{X}]$. We can therefore speak about the Frobenius fixed points $R^{F}$, ${X}^{F}$, and $[{X}]^{F}$. The first two are finite sets while the last one is a finite groupoid.
3.3.1. Points of the representation variety
By definition,
@@ 3.3.1. Points of the representation variety
Henceforth, we assume that the stabiliser in $G$ of each element in $C_{i}$ is connected. It then follows that $C_{i}^{F}$ is a unique $G^{F}$-conjugacy class, cf. [GM20, §2.6.1]. Thus, it makes sense to evaluate irreducible complex characters of $G^{F}$ on $C_{i}^{F}$. The following result goes back to the work of Frobenius; see e.g. [HLRV11, §3] for a proof:
Theorem 14 (Frobenius Mass Formula).
@@ 3.3.3. Points of character variety
We now assume that the tuple $(C_{1},\ldots,C_{n})$ is generic. We saw above that ${X}$ is a coarse moduli space for the smooth Deligne-Mumford stack $[{X}]$. This implies that $[{X}]$ and $X$ have the same number of points over finite fields, cf. [Behrend91, §2] or [BogaartEdixhoven]. Thus, we have
@@ 3.4. Additive character varieties
Let $\mathcal{O}_{1},\ldots,\mathcal{O}_{n}$ be $F$-stable $G$-orbits in $\mathfrak{g}$. Let $A$ denote the corresponding additive representation variety. Then
@@
Let $1_{\mathcal{O}_{i}}^{G}$ denote the characteristic function of the $G^{F}$-orbit $\mathcal{O}_{i}^{F}$. Let $\mathcal{F}$ denote the Fourier transform on the space of $G^{F}$-invariant functions on $\mathfrak{g}^{F}$ (see §5.1 for the definition). For each $x\in\mathfrak{g}$, let $\mathfrak{g}_{x}$ denote its centraliser in $\mathfrak{g}$. Assume that the stabiliser in $G$ of every element of $\mathcal{O}_{i}$ is connected. Thus, $\mathcal{O}_{i}^{F}\subset\mathfrak{g}^{F}$ is a unique $G^{F}$-orbit. Then we have (cf. [HLRV11, §3]):
@@
If $\mathcal{O}=(\mathcal{O}_{1},\ldots,\mathcal{O}_{n})$ is generic, then

### v2

@@ 2.1.2. Dual root systems
For each root $\alpha\in\Phi$, define the coroot
$\check{\alpha}:=\frac{2}{(\alpha,\alpha)}\alpha\in V.$
The set of coroots forms the dual root system $\check{\Phi}\subset V$. If $\Psi$ is a subsystem of $\Phi$, then $\check{\Psi}$ is a subsystem of $\check{\Phi}$. However, if $\Psi$ is closed in $\Phi$, then $\check{\Psi}$ need not be closed in $\check{\Phi}$. For instance, $C_{n}\times C_{n}$ is a closed subsystem of $C_{2n}$ but if $n>1$, then $B_{n}\times B_{n}$ is not closed in $B_{2n}$. Indeed, $B_{n}\times B_{n}$ does not arise in the Borel–de Siebenthal algorithm applied to $B_{2n}$.
@@ 2.1.3. Levi subsystems
A Levi subsystem (aka parabolic subsystem) of $\Phi$ is a subsystem of the form $\Phi\cap E$ where $E\subseteq V$ is a subspace. A Levi subsystem of $\Phi$ is closed in $\Phi$. Moreover, every Levi subsystem of $\Phi$ is of the form $w\cdot\langle S\rangle$, where $S$ is a subset of $\Delta$ and $w\in W$. Note that $\Psi$ is a Levi subsystem of $\Phi$ if and only if $\check{\Psi}$ is a Levi subsystem of $\check{\Phi}$.
@@ 2.1.4. Isolated subsystems
A subset $S\subseteq\Phi$ is called isolated in $\Phi$ if $S$ is not contained in a proper Levi subsystem of $\Phi$. The following are equivalent:
@@ 2.1.5. Pseudo-Levi subsystems
Assume $\Phi$ is irreducible. Let $\theta$ be the highest root of $\Phi$, and
$\widetilde{\Delta}:=\Delta\sqcup\{-\theta\}.$
A pseudo-Levi subsystem is a subsystem of $\Phi$ of the form $w\cdot\langle S\rangle$, where $w\in W$ and $S$ is a subset of $\widetilde{\Delta}$. More generally, if $\Phi=\Phi_{1}\sqcup\cdots\sqcup\Phi_{r}$ with $\Phi_{i}$ irreducible, then a subsystem $\Psi\subseteq\Phi$ is said to be a pseudo-Levi if $\Psi_{i}:=\Psi\cap\Phi_{i}$ is a pseudo-Levi inside $\Phi_{i}$ for all $i$. It is clear that every Levi is a pseudo-Levi. The converse is true in type $A$ but false otherwise. For instance, $C_{2}$ has a pseudo-Levi $A_{1}\times A_{1}$ which is not a Levi.
@@ 2.1.6. Isolated pseudo-Levis
Let $\Psi$ be a pseudo-Levi subsystem of an irreducible root system $\Phi$. Then the following are equivalent:
@@
    $\Psi=w\cdot\langle S\rangle$ where $S\subset\tilde{\Delta}$ has the same size as $\Delta$.
@@ 2.2.3. Parabolic subgroups
A connected subgroup $P\subseteq G$ is called parabolic if the quotient $G/P$ is proper. Every parabolic subgroup has a Levi decomposition $P=L\ltimes N$, where $N$ is the maximal closed connected normal unipotent subgroup of $P$ and $L$ is a Levi subgroup of $G$. Every Levi subgroup of $G$ arises in this manner.
2.2.4.
In what follows, we will be using the following facts:
@@ 2.2.5. Pseudo-Levi subgroups
This is a pseudo-Levi subsystem of $\Phi$. Moreover, every pseudo-Levi subsystem arises in this manner [Deriziotis]. By a theorem of Steinberg, $\pi_{0}(G_{x})$ is isomorphic to a subgroup of $\pi_{1}([G,G])$; thus, if $[G,G]$ is simply connected, then $G_{x}$ is connected [Steinberg75, §2].
@@ 2.2.7. Quasi-isolated elements
A semisimple element $x\in G$ is called quasi-isolated if $G_{x}$ is not in a proper Levi subgroup of $G$. It is called isolated if $G_{x}^{\circ}$ is not in any proper Levi subgroup of $G$. Thus, $x\in T$ is isolated if and only if $\Phi_{x}$ is an isolated (pseudo-Levi) subsystem of $\Phi$. If $x$ is quasi-isolated, then $G_{x}$ is an irreducible subgroup of $G$; i.e., it does not belong to a proper parabolic. By Schur’s lemma (cf. [Sikora, Proposition 15]) $Z(G_{x})/Z(G)$ is finite.
@@ 3. Generic character varieties
Let $G$ be a connected reductive group over an algebraically closed field $k$. In this section, we define the notion generic tuples of conjugacy classes of $G$ and study the corresponding character varieties. Most of this material appears in [HLRV11] for $\mathrm{GL}_{n}$ and [Boalch14] for general $G$ over complex numbers. We work with reductive groups over arbitrary fields.
@@ 3.1.1. Maximal parabolics
Clearly it is enough to check the above condition on maximal (proper) parabolic subgroups. This implies that if $G=\mathrm{GL}_{n}$, then being generic is equivalent to requiring that whenever a proper non-zero subspace $V\subset k^{n}$ is preserved by $x_{i}\in C_{i}$, $i=1,\ldots,n$, we have $\prod_{i=1}^{n}\mathrm{det}(x_{i}|_{V})\neq 1$. Thus, we recover the notion of generic tuples of [HLRV11, §2].
@@ 3.1.2. Standard parabolics
It is enough to check the genericity condition for standard (proper) parabolic subgroups. (By definition, standard parabolic subgroups are those containing a chosen Borel subgroup $B\subset G$.) Indeed, suppose $P$ is an arbitrary parabolic and choose $g\in G$ such that the conjugate $g\cdot P$ is standard. Then
$x_{i}\in P\cap C_{i}\qquad\iff\qquad g\cdot x_{i}\in g\cdot P\cap C_{i}$
and
$\prod_{i=1}^{n}x_{i}\in[P,P]\qquad\iff\qquad g\cdot\prod_{i=1}^{n}x_{i}\in g\cdot[P,P]=[g\cdot P,g\cdot P].$
@@ 3.1.3. Semisimple part
For each conjugacy class $X$, write $X^{s}$ for the corresponding semisimple class, obtained by taking the semisimple parts of the elements of $X$. Then one can show that $C=(C_{1},\ldots,C_{n})$ is generic if and only if $C^{s}=(C_{1}^{s},\ldots,C_{n}^{s})$ is generic. This is the reductive analogue of the fact that in $\mathrm{GL}_{n}$, being generic depends only on the eigenvalues.
@@ 3.1.5. Generic elements
We now explain that deciding whether a tuple of (semisimple) conjugacy classes is generic involves checking finitely many conditions. To this end, we need a definition.
@@
A tuple $S=(S_{1},\ldots,S_{n})\in T^{n}$ is said to be generic if for every $(w_{1},\ldots,w_{n})\in W^{n}$ and every proper standard Levi subgroup $L\subset G$, we have
$\prod_{i=1}^{n}w_{i}\cdot S_{i}\notin T\cap[L,L].$
Let $C_{i}:=G\cdot S_{i}$ be the conjugacy class corresponding to $S_{i}$.
@@
It is clear that if $C$ is generic then so is $S$. For the converse, let $x=(x_{1},\ldots,x_{n})\in C_{1}\times\ldots\times C_{n}$. Let $L\subset G$ be a proper standard Levi subgroup containing $x_{1},\ldots,x_{n}$. We need to show $\prod_{i=1}^{n}x_{i}\notin[L,L]$. We have $x_{i}=l_{i}\cdot t_{i}$ for some $l_{i}\in L$ and $t_{i}\in T$. Now $t_{i}$ and $S_{i}$ are elements of $T$ which are $G$-conjugate; thus, they must actually be conjugate under $W$; i.e., there exists $w_{i}\in W$ such that $t_{i}=w_{i}\cdot S_{i}$, cf. proof of Proposition 3.7.1 in [Carter93]. By assumption, $t_{1}\cdots t_{n}\notin[L,L]$ which implies that $x_{1}\cdots x_{n}\notin[L,L]$. Thus, $C$ is generic. ∎
@@ 3.1.6. Existence
The above lemma shows that generic conjugacy classes exist. Indeed, if $L$ is a proper standard levi subgroup, then $T\cap[L,L]$ is a proper closed subvariety of $T$. (At the level of Lie algebras, $\mathfrak{t}=\mathfrak{t}\cap[\mathfrak{l},\mathfrak{l}]\oplus Z(\mathfrak{l})$ for $\mathfrak{t}=\mathrm{Lie}(T)$ and the centre $Z(\mathfrak{l})$ is nontrivial if $\mathfrak{l}$ is a proper Levi subalgebra of $\mathfrak{g}$.) Since there are only finitely many standard Levis $L$ and finitely many $w$’s, we conclude that the generic locus of $T^{n}$ is open dense.
@@
Let $\mathbf{T}_{n}\subseteq T^{n}$ denote the subvariety consisting of those tuples $(S_{1},\ldots,S_{n})\in T^{n}$ satisfying $\prod_{i=1}^{n}S_{i}\in[G,G]$. Then the above argument also shows that the generic locus of $\mathbf{T}_{n}$ is open dense. Since the (strongly) regular locus is also open dense in $T^{n}$, it follows that conjugacy classes satisfying Assumption 1 exist.
@@ 3.2. Generic character varieties
Let ${r=(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in R}$ and suppose $P$ is a parabolic subgroup of $G$ containing all $a_{i},b_{i},$ and $c_{i}$’s. Then, $[P,P]$ contains $[a_{1},b_{1}]\cdots[a_{g},b_{g}]$; thus, $[P,P]$ also contains $\prod_{i=1}^{n}c_{i}$. As $C$ is generic, it follows that $P=G$. Thus, $r$ is irreducible and Schur’s lemma (cf. [Sikora, §4]) implies that $\mathrm{Stab}_{G/Z}(r)$ is finite. Note that this stabiliser is the same as the centraliser $C_{G}(H)$, where $H:=\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle\subseteq G$. By Theorem 10, this group is smooth. Since we already have shown that it is finite, it follows that it is finite and étale.
Now stability follows from the fact that whenever a connected reductive group acts on an affine scheme of finite type with finite stabilisers, then every point is stable, cf. [Mumford, Proposition 0.8]. (Alternatively, one can use the Hilbert–Mumford criterion as in [Boalch14].) The statements about $R$ follow from the fact that $1$ is a regular value for $M$, see [KNP23, §2] for details. ∎
@@
Next, let ${X}:=R/\!\!/G$ and $[{X}]:=[R/(G/Z)]$ denote, respectively, the (generic) character variety and character stack associated to $(\Gamma_{g,n},G,C)$. The previous theorem implies:
@@ 3.3. Frobenius Mass Formula
We now work over $k=\overline{\mathbb{F}_{q}}$ and assume that $G$ is equipped with an untwisted Frobenius automorphism $F:G\rightarrow G$. We also assume that the conjugacy classes $C=(C_{1},\ldots,C_{n})$ are $F$-stable. Then the action of the Frobenius $F$ on $G$ and $C_{i}$’s, gives an action of $F$ on $R$, ${X}$ and $[{X}]$. We can therefore speak about the Frobenius fixed points $R^{F}$, ${X}^{F}$, and $[{X}]^{F}$. The first two are finite sets while the last one is a finite groupoid. By definition,
@@
3.3.1. Stable conjugacy
Definition 17.
A conjugacy class $C\subseteq G$ is called stable if the stabiliser in $G$ of each element of $C$ is connected.
If $C$ is stable, then $C^{F}$ is a unique $G^{F}$-conjugacy class, cf. [GM20, §2.7.1]. Thus, it makes sense to evaluate irreducible complex characters of $G^{F}$ on $C_{i}^{F}$.
3.3.2. Frobenius Mass Formula
The following result goes back to the work of Frobenius; see e.g. [HLRV11, §3] for a proof:
Theorem 18.
If $C_{i}$’s are stable, then
@@ 3.3.4. Points of character variety
We now assume that the tuple $(C_{1},\ldots,C_{n})$ is generic. We saw above that ${X}$ is a coarse moduli space for the smooth Deligne-Mumford stack $[{X}]$. This implies that $[{X}]$ and $X$ have the same number of points over finite fields, cf. [Behrend91, §2] or [BogaartEdixhoven]. Thus, we conclude:
Corollary 19.
If $C_{i}$’s are generic and stable, then
@@ 3.4. Additive character varieties
Let $\mathcal{O}_{1},\ldots,\mathcal{O}_{n}$ be $F$-stable adjoint $G$-orbits in $\mathfrak{g}$. Let $A$ denote the corresponding additive representation variety. Then
@@
Let $1_{\mathcal{O}_{i}}^{G}$ denote the characteristic function of the $G^{F}$-orbit $\mathcal{O}_{i}^{F}$. Let $\mathcal{F}$ denote the Fourier transform on the space of $G^{F}$-invariant functions on $\mathfrak{g}^{F}$ (see §6.1 for the definition). For each $x\in\mathfrak{g}$, let $\mathfrak{g}_{x}$ denote its centraliser in $\mathfrak{g}$. Then we have (cf. [HLRV11, §3]):
@@
If $\mathcal{O}_{i}$ are generic and stable, then